OpenStudy (anonymous):

What is the solution to the proportion? 14/16=21/x A-x=10.7 B-x=18.4 C-x=24 D-x=32

OpenStudy (anonymous):

@crabby_patty

OpenStudy (anonymous):

x = 24

OpenStudy (anonymous):

What is the solution to the proportion? 3/m = 12/15 A. m = 2.4 B. m = 3.75 C. m = 4.5 D. m = 4.75

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Fifteen note cards cost $4.80. How much does it cost to buy 8 note cards? A. $25.00 B. $9.00 C. $3.20 D. $2.56

OpenStudy (anonymous):

D ) 2.56

OpenStudy (anonymous):

A pastry recipe calls for 3 cups of flour to make 8 servings. How many cups of flour are needed to make 6 servings? A. 2.25 cups B. 2.5 cups C. 4.5 cups D. 16 cups

OpenStudy (anonymous):

A

OpenStudy (anonymous):

anything else??

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome...anytime